OpenStudy (anonymous):

Write the function in slope intercept form. Then graph the function. -y = -8x (NO IDEA WHAT I'M DOING!! Please help!)

OpenStudy (fifciol):

multiply both sides by -1

OpenStudy (anonymous):

-y=8x ?

OpenStudy (fifciol):

no, you've only multiplied right side you have to multiply left side too

OpenStudy (anonymous):

How?

OpenStudy (fifciol):

\[-y=-8x/*(-1)\] \[-y*(-1)=-8x*(-1)\]

OpenStudy (fifciol):

in other words you have to find y, that's why we multiply both sides by -1

OpenStudy (fifciol):

so what result you have?

OpenStudy (fifciol):

It's y=8x right?

OpenStudy (anonymous):

I guess.. This is confusing

OpenStudy (fifciol):

now you have to graph this function y=8x. You take some two randomly chosen x's (for example x=0 and x=1) What values of y's will you have?(just substitute x's to equation)

OpenStudy (fifciol):

for x =0 you get y=8*0=0. You have your first point (0,0) then for x=1 you get 8*1=8. You have your second point (1,8)

OpenStudy (fifciol):